OpenStudy (nuccioreggie):

Convert 1 over 8 to a decimal and tell whether it terminates or repeats. 0.125, a number that terminates 1.8, a number that terminates 1.8 with a repeating bar over 8., a number that repeats 0.875 with a repeating bar over 875., a number that repeats

OpenStudy (bloomlocke367):

do you have a calculator?

OpenStudy (anonymous):

0.125, a number that terminates.

OpenStudy (anonymous):

fan and meadl @nuccioreggie and @ me for more help

OpenStudy (nuccioreggie):

@ganeshie8

ganeshie8 (ganeshie8):

you want to find the decimal form of 1/8

ganeshie8 (ganeshie8):

1/2 = one half = 0.5 1/4 = one quarter = 0.25 1/8 = ?

OpenStudy (nuccioreggie):

is it A

OpenStudy (nuccioreggie):

?

OpenStudy (nuccioreggie):

1.8

ganeshie8 (ganeshie8):

Yep! the number terminates because 8 = 2^3 and 2 is a factor of 10

OpenStudy (nuccioreggie):

lol so its C

OpenStudy (nuccioreggie):

@ganeshie8

OpenStudy (nuccioreggie):

i ment B